OpenStudy (anonymous):

In right ∆PQR, P and Q are acute angles, and cos P is 3/5. What is sin Q? Answers: 4/5 3/5 2/5 3/4

Nnesha (nnesha):

\[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]